Find all negative odd integers that satisfy the following inequality
-3x+1=<17

Answer by josmiceli(6784)

-3x + 1 =< 17
-3x =< 16
x => -5.33
The negative odd integers that satisfy this are -5, -3, -1
check
-3(-5) + 1 =< 17
15 + 1 =< 17
TRUE
-3(-3) + 1 =< 17
9 + 1 =< 17
TRUE
-3(-1) + 1 =< 17
3 +1 =< 17
TRUE
note that -7 or less does not satisfy equation